# Scene Flow as a Partial Differential Equation

### v1

@@ 5.2 What contributes to EulerFlow’s state-of-the-art performance?
We find that EulerFlow’s lower mean Dynamic EPE can be attributed to better performance on smaller objects. On Argoverse 2, compared to Flow4D, EulerFlow’s improves on WHEELED VRU (Figure 6(d)), a small, rare, fast moving class. Compared to ICP Flow, EulerFlow’s improves on all classes (at least halving the error on every class!), with the largest improvements coming from the smaller and harder to detect objects PEDESTRAIN and WHEELED VRU (Figures 6(c)6(d)). On Waymo Open, the same story holds; the most dramatic performance improvements come from the small object classes of CYCLIST and PEDESTRIAN (Figure 7).
These results are consistent with our qualitative visualizations. Figure 12 shows EulerFlow is able to cleanly extract the motion of a bird flying past the ego vehicle. Euler integration using EulerFlow’s PDE, starting at the bird’s takeoff position and ending when it loses lidar returns, produces emergent 3D point tracking behavior on the bird through its trajectory (Figure 8), further demonstrating the quality of EulerFlow’s model of the scene’s motion.
@@ 5.2.1 How does observation sequence length impact EulerFlow?
As we discuss in Section 3, EulerFlow benefits from constructive interference from PDE estimation over many observations. Does this sufficiently explain EulerFlow’s performance? Figure 9 shows the performance of EulerFlow at length 5, 20, 50, and full sequence (roughly 160 frames) compared to NSFP and NTP at length 20. EulerFlow sees clear continual improvements as the number of frames increases without signs of saturation. However, sequence length alone does not explain EulerFlow’s performance; even at the same sequence length of 20, EulerFlow demonstrates significantly better performance than NTP.
@@ 5.2.3 How does the capacity of the neural prior impact EulerFlow?
Li et al. ablate the capacity of NSFP’s neural prior to characterize underfitting and overfitting to optimization objective noise, ultimately selecting a depth of 8. EulerFlow’s neural prior is structured similarly; however, NSFP is fitting a single snapshot in time, while EulerFlow is fitting an entire PDE over significant time intervals. Intuitively, one would expect that full sequence modeling would benefit from greater network capacity.
To evaluate this, we perform a sweep of EulerFlow’s network depth on the Argoverse 2 validation split (Figure 11). While EulerFlow with NSFP’s default of depth 8 performs well on our Argoverse 2 evaluations (0.1% worse than the supervised state-of-the-art Flow4D), we see that performance improves as the neural prior’s depth increases until depth 18 (indicating underfitting), where we then start to see degradation (indicating overfitting to surrogate objective noise). Based on these results, as we discuss in Section 5.1, our submission to the Argoverse 2 2024 Scene Flow Challenge leaderboard uses a depth 18 neural prior (Figure 4).
Figure 11: Mean Dynamic Normalized EPE of EulerFlow on the Argoverse 2 val split for different neural prior capacities. Shallow networks underfit the PDE, while deeper networks overfit to noise in the optimization objectives.
@@ 5.3 Beyond Autonomous Vehicles
Due to a dearth of real-world, labeled scene flow data, prior scene flow work on real data overwhelmingly evaluates on autonomous vehicle datasets (Dewan et al., 2016; Li et al., 2021b; Jund et al., 2021; Li et al., 2023; Chodosh et al., 2023; Liu et al., 2024; Vedder et al., 2024; Khatri et al., 2024); consequently, motion understanding in other important domains like tabletop manipulation has been neglected. To showcase EulerFlow’s out-of-the-box flexibility and generalizability, we visualize EulerFlow on several dynamic tabletop scenes we collected using the ORBBEC Astra, a low cost depth camera commonly used in robotics (Figure 13). For viewing ease, we paint our point clouds with color; however, RGB information is not provided to EulerFlow during optimization. Furthermore, while EulerFlow only reasons about point clouds, we show that it can leverage video monocular depth estimators to describe RGB-only scene flow in Supp. A.2. Interactive visuals are available at vedder.io/eulerflow.
Figure 13: Visualizations of EulerFlow’s emergent 3D point tracking behavior that demonstrate the quality of its PDE estimate. Row 1 depicts tracking a tomato placed in the sink by a human hand; note the point does not move despite the hand grasping the tomato. Row 2 depicts tracking of painters tape rolling off a table; EulerFlow is able to estimate its trajectory even after it disappears out of frame. Row 3 depicts tracking of the motion of a jack commonly used in tabletop manipulation experiments (Venkatesh et al., 2023). Row 4 depicts tracking of a tennis ball taped to a flexible rod. All tracks are produced by Euler integration through the estimated PDE from the initial conditions shown in the left column. Note that point clouds are shown in color for visualization purposes only; RGB is not used during optimization.
@@ 6 Conclusion
By reframing scene flow as fitting a PDE over positions for a full sequence of observations, we are able to construct EulerFlow, a simple unsupervised scene flow method that achieves state-of-the-art performance on the Argoverse 2 2024 Scene Flow Challenge and Waymo Scene Flow benchmark, where it beats all prior art, supervised or unsupervised. EulerFlow is able to describe motion on small, fast moving, out of distribution objects unable to be captured by prior art, suggesting that it makes good on the promises of scene flow as a powerful primitive for understanding the dynamic world. It also exhibits other emergent capabilities, like basic 3D point tracking behavior.
We believe that this PDE formulation has implications for scene flow at large, including beyond test-time optimization methods; the power of multi-step Euler integration may translate to feedforward network training. Future work should explore feedforward models that perform autoregressive rollouts or directly learn to estimate multiple steps into the future.
@@ 6.1 Limitations and Future Work
EulerFlow is expensive to optimize. With our implementation, optimizing EulerFlow for a single Argoverse 2 sequence takes 24 hours on one NVIDIA V100 16GB GPU, putting it on par with the original NeRF paper’s computation expense (Mildenhall et al., 2021). However, like with NeRF, we believe further optimization, algorithmic, and engineering innovation can significantly improve runtime.
@@ A.1 How does the choice of learnable function class and design of encodings impact EulerFlow?
One of Li et al.’s core theoretical contributions demonstrates that NSFP’s ReLU MLP is a good prior for scene flow because it represents a smooth learnable function class, and scene flow is often locally smooth with respect to input position. However, unlike NSFP, EulerFlow is fitting flow over a full PDE; while it seems reasonable to assume that this PDE is typically also locally smooth, cases like adjacent cars moving rapidly in opposite directions may benefit from the ability to model higher frequency, less locally smooth functions. To test this hypothesis, we ablate EulerFlow by replacing its normalized time with higher frequency sinusoidal time embeddings (mirroring Wang et al.’s proposed time embedding for NTP), as well as try other popular non-linearities like SinC (Ramasinghe et al., 2024) and Gaussian (Chng et al., 2022) from the coordinate network literature. Figure 14 features negative results on these ablations across the board; Gaussians were unable to converge due the extremely high frequency representation triggering early stopping, while the use of SinC and higher frequency time embeddings both resulted in worse overall performance, indicating that Li et al.’s smooth function prior does indeed seem appropriate for EulerFlow’s neural prior.
@@ B.2 Why didn’t you use a Neural ODE or a Liquid Neural Network?
Neural ODEs (Chen et al., 2018) take variable size and number of steps in latent space to do inference; imagine a ResNet that can use an ODE solver to dynamically scale the impact of the residual block, as well as decide the number of residual blocks. They are not a function class specially designed to fit a PDE well. Similar to Neural ODEs, Liquid Neural Networks (Hasani et al., 2021) focus on the same class of problems and are similarly not applicable.
@@ B.3 Why didn’t you do experiments on FlyingThings3D / <simulated dataset>?
Most popular synthetic datasets do not contain long observation sequences (Mayer et al., 2016; Butler et al., 2012), but instead include standalone frame pairs. Our method leverages the long sequence of observations to refine our neural estimate of the true PDE. Indeed, on two frames, EulerFlow collapses to NSFP.
@@
Instead, we want to turn our attention to the sort of workloads that do clearly matter — describing motion in domains like manipulation or autonomous vehicles, where it seems clear that scene flow, if solved, will serve as powerful primitive for downstream systems. This is why we performed qualitative experiments on the tabletop data we collected ourselves; to our knowledge, no real-world dynamic datasets of this nature exist with ground truth annotations, but we want to emphasize that EulerFlow works in such domains, and consequently EulerFlow and other Scene Flow via PDE-based methods can be used as a primitive in these real world domains.

### v2

@@ 5.2 What contributes to EulerFlow’s state-of-the-art performance?
We find that EulerFlow’s lower mean Dynamic EPE can be attributed to better performance on smaller objects. On Argoverse 2, compared to Flow4D, EulerFlow’s improves on WHEELED VRU (Figure 7(d)), a small, rare, fast moving class. Compared to ICP Flow, EulerFlow’s improves on all classes (at least halving the error on every class!), with the largest improvements coming from the smaller and harder to detect objects PEDESTRAIN and WHEELED VRU (Figures 7(c)7(d)). On Waymo Open, the same story holds; the most dramatic performance improvements come from the small object classes of CYCLIST and PEDESTRIAN (Figure 8).
These results are consistent with our qualitative visualizations. Figure 13 shows EulerFlow is able to cleanly extract the motion of a bird flying past the ego vehicle. Euler integration using EulerFlow’s ODE, starting at the bird’s takeoff position and ending when it loses lidar returns, produces emergent 3D point tracking behavior on the bird through its trajectory (Figure 9), further demonstrating the quality of EulerFlow’s model of the scene’s motion.
@@ 5.2.1 How does observation sequence length impact EulerFlow?
As we discuss in Section 3, EulerFlow benefits from constructive interference from ODE estimation over many observations. Does this sufficiently explain EulerFlow’s performance? Figure 10 shows the performance of EulerFlow at length 5, 20, 50, and full sequence (roughly 160 frames) compared to NSFP and NTP at length 20. EulerFlow sees clear continual improvements as the number of frames increases without signs of saturation. However, sequence length alone does not explain EulerFlow’s performance; even at the same sequence length of 20, EulerFlow demonstrates significantly better performance than NTP.
@@ 5.2.3 How does the capacity of the neural prior impact EulerFlow?
Li et al. ablate the capacity of NSFP’s neural prior to characterize underfitting and overfitting to optimization objective noise, ultimately selecting a depth of 8. EulerFlow’s neural prior is structured similarly; however, NSFP is fitting a single snapshot in time, while EulerFlow is fitting an entire ODE over significant time intervals. Intuitively, one would expect that full sequence modeling would benefit from greater network capacity.
To evaluate this, we perform a sweep of EulerFlow’s network depth on the Argoverse 2 validation split (Figure 12). While EulerFlow with NSFP’s default of depth 8 performs well on our Argoverse 2 evaluations (0.1% worse than the supervised state-of-the-art Flow4D), we see that performance improves as the neural prior’s depth increases until depth 18 (indicating underfitting), where we start to see degradation (indicating overfitting to noise). Based on these results our Argoverse 2 2024 Scene Flow Challenge leaderboard submission uses a depth 18 neural prior (Figure 5).
Figure 12: Mean Dynamic Normalized EPE of EulerFlow on the Argoverse 2 val split for different neural prior capacities. Shallow networks underfit the ODE, while deeper networks overfit to noise in the optimization objectives.
@@ 5.3 Beyond Autonomous Vehicles
Due to a dearth of real-world, labeled scene flow data, prior scene flow work on real data overwhelmingly evaluates on autonomous vehicle datasets (Dewan et al., 2016; Li et al., 2021b; Jund et al., 2021; Li et al., 2023; Chodosh et al., 2023; Liu et al., 2024; Vedder et al., 2024; Khatri et al., 2024); consequently, motion understanding in other important domains like tabletop manipulation has been neglected. To showcase EulerFlow’s out-of-the-box flexibility and generalizability, we visualize EulerFlow on several dynamic tabletop scenes we collected using the ORBBEC Astra, a low cost depth camera commonly used in robotics (Figure 14). For viewing ease, we paint our point clouds with color; however, RGB information is not provided to EulerFlow during optimization. While EulerFlow only reasons about point clouds, it can leverage video mono depth estimates to describe RGB-only scene flow (Appendix A.2). Interactive visuals are available at vedder.io/eulerflow.
Figure 14: Visualizations of EulerFlow’s emergent 3D point tracking behavior that demonstrate the quality of its ODE estimate. Row 1 depicts tracking a tomato placed in the sink by a human hand; note the point does not move despite the hand grasping the tomato. Row 2 depicts tracking of painters tape rolling off a table; EulerFlow is able to estimate its trajectory even after it disappears out of frame. Row 3 depicts tracking of the motion of a jack commonly used in tabletop manipulation experiments (Venkatesh et al., 2023). Row 4 depicts tracking of a tennis ball taped to a flexible rod. All tracks are produced by Euler integration through the estimated ODE from the initial conditions shown in the left column. Note that point clouds are shown in color for visualization purposes only.
@@ 6 Conclusion
By reframing scene flow as fitting an ODE over positions for a full sequence of observations, we are able to construct EulerFlow, a simple unsupervised scene flow method that achieves state-of-the-art performance on the Argoverse 2 2024 Scene Flow Challenge and Waymo Scene Flow benchmark, where it beats all prior art, supervised or unsupervised. EulerFlow is able to describe motion on small, fast moving, out of distribution objects unable to be captured by prior art, suggesting that it makes good on the promises of scene flow as a powerful primitive for understanding the dynamic world. It also exhibits other emergent capabilities, like basic 3D point tracking behavior.
We believe that this ODE formulation has implications for scene flow at large, including beyond test-time optimization methods; the power of multi-step Euler integration may translate to feedforward network training. Future work should explore feedforward models that perform autoregressive rollouts or directly learn to estimate multiple steps into the future.
@@ 6.1 Limitations and Future Work
EulerFlow is expensive to optimize. With our implementation, optimizing EulerFlow for a single Argoverse 2 sequence takes 24 hours on one NVIDIA V100 16GB GPU, putting it on par with the original NeRF paper’s computation expense (Mildenhall et al., 2021). However, like with NeRF, we believe algorithmic, optimization, and engineering improvements can significantly reduce runtime.
@@ A.1 How does the choice of learnable function class and design of encodings impact EulerFlow?
One of Li et al.’s core theoretical contributions demonstrates that NSFP’s ReLU MLP is a good prior for scene flow because it represents a smooth learnable function class, and scene flow is often locally smooth with respect to input position. However, unlike NSFP, EulerFlow is fitting flow over a full ODE; while it seems reasonable to assume that this ODE is typically also locally smooth, cases like adjacent cars moving rapidly in opposite directions may benefit from the ability to model higher frequency, less locally smooth functions. To test this hypothesis, we ablate EulerFlow by replacing its normalized time with higher frequency sinusoidal time embeddings (mirroring Wang et al.’s proposed time embedding for NTP), as well as try other popular non-linearities like SinC (Ramasinghe et al., 2024) and Gaussian (Chng et al., 2022) from the coordinate network literature. Figure 15 features negative results on these ablations across the board; Gaussians were unable to converge due the extremely high frequency representation triggering early stopping, while the use of SinC and higher frequency time embeddings both resulted in worse overall performance, indicating that Li et al.’s smooth function prior does indeed seem appropriate for EulerFlow’s neural prior.
@@ B.2 Why didn’t you use a Neural ODE or a Liquid Neural Network?
Neural ODEs (Chen et al., 2018) take variable size and number of steps in latent space to do inference; imagine a ResNet that can use an ODE solver to dynamically scale the impact of the residual block, as well as decide the number of residual blocks. They are not a function class specially designed to fit derivative estimates well. Similar to Neural ODEs, Liquid Neural Networks (Hasani et al., 2021) focus on the same class of problems and are similarly not applicable.
@@ B.3 Why didn’t you do experiments on FlyingThings3D / <simulated dataset>?
Most popular synthetic datasets do not contain long observation sequences (Mayer et al., 2016; Butler et al., 2012), but instead include standalone frame pairs. Our method leverages the long sequence of observations to refine our neural estimate of the true ODE. Indeed, on two frames, EulerFlow collapses to NSFP.
@@
Instead, we want to turn our attention to the sort of workloads that do clearly matter — describing motion in domains like manipulation or autonomous vehicles, where it seems clear that scene flow, if solved, will serve as powerful primitive for downstream systems. This is why we performed qualitative experiments on the tabletop data we collected ourselves; to our knowledge, no real-world dynamic datasets of this nature exist with ground truth annotations, but we want to emphasize that EulerFlow works in such domains, and consequently EulerFlow and other Scene Flow via ODE-based methods can be used as a primitive in these real world domains.
Appendix C EulerFlow implementation details
Our neural prior $\theta$ is a straightforward extension to NSFP’s coordinate network prior [^6] Hyperparameters (e.g. filter width of 128) of NSFP’s prior are kept fixed, except for depth (Section 5.2.3).; however, instead of taking a 3D space vector (positions $X,Y,Z\in\mathbb{R}$) as input, we encode a 5D space-time-direction vector: positions $X,Y,Z,\in\mathbb{R}$, sequence normalized time $t\in[-1,1]$ (i.e. the point cloud time scaled to this range), and direction $d\in\{\texttt{BWD}=-1,\texttt{FWD}=1\}$. This simple encoding scheme enables description of arbitrary regions of the ODE, allowing for the ODE to be queried at frequencies different from the sensor frame rate. Euler integration enables simple implementation of multi-step forward, backward, and cyclic consistency losses without extra bells and whistles. For efficiency, we use Euler integration with $\Delta{}t$ set as the time between observations for our ODE solver, enabling support for arbitrary sensor frame rates, and set the cycle consistency balancing term $\alpha=0.01$ and optimization window $W=3$ for all experiments.
Appendix D EulerFlow’s ODE Derivation
D.1 Formulating the ODE
Given a (possibly moving) particle in some canonical frame (i.e. time 00), we define a function $L(x_{0},y_{0},z_{0},t)$ that can describe its location at an arbitrary future time $t$, i.e. a Lagrangian description of motion (Figure 4).
$L(x_{0},y_{0},z_{0},t)=x_{t},y_{t},z_{t}$ (4)
For notational clarity to access $x_{t},y_{t},z_{t}$ individually, we can define
$L_{x}(x_{0},y_{0},z_{0},t)$ $=x_{t}$ (5)
$L_{y}(x_{0},y_{0},z_{0},t)$ $=y_{t}$ (6)
$L_{z}(x_{0},y_{0},z_{0},t)$ $=z_{t}$ (7)
Similarly, we can define $F(x_{t},y_{t},z_{t},t)$ to describe the instantaneous velocity of a point $x_{t},y_{t},z_{t}$ at some arbitrary time $t$, i.e. a Eulerian description of motion (Figure 4).
$\frac{dL(x_{0},y_{0},z_{0},t)}{dt}=\frac{dL}{dt}=\left(\frac{dL_{x}}{dt},\frac{dL_{y}}{dt},\frac{dL_{z}}{dt}\right)=F(x_{t},y_{t},z_{t},t)$ (8)
$F$ is defined in terms of the total derivative of $L$ with respect to $t$, as $x_{0},y_{0},z_{0}$ are initial conditions that do not vary with time (i.e. $\frac{dL}{dt}=\frac{\partial L}{\partial t}+\frac{\partial L}{\partial x_{0}}\frac{dx_{0}}{dt}+\frac{\partial L}{\partial y_{0}}\frac{dy_{0}}{dt}+\frac{\partial L}{\partial z_{0}}\frac{dz_{0}}{dt}=\frac{\partial L}{\partial t}$, as $\frac{dx_{0}}{dt}=\frac{dy_{0}}{dt}=\frac{dz_{0}}{dt}=0$). We can exactly define $L$ recursively in terms of the initial conditions and $F$, i.e.
[…]
or, more compactly,
[…]
Our function $L$ can thus be defined as a multi-dimensional ODE in terms of $F$ with initial conditions $x_{0},y_{0},z_{0}$.
D.2 Arbitrary start and end times from the Eulerian formulation
In the above derivation, $L$ requires that a moving point be defined in terms of a canonical frame defined at time 00, as is common in the deformation in reconstruction literature. However, the Eulerian formulation has no such requirement, allowing us to select arbitrary start and end times across different point queries. To showcase this, we can query $F$ to extract the trajectory of a particle at $t$ across the range $[t,t^{\prime}]$ starting at $x_{t},y_{t},z_{t}$ simply by changing the range of the integral in Equation 10, i.e.
[…]
While $E$ and $L$ appear similar on their face, $E$ is strictly more flexible than $L$. In principle you could choose to redefine $L$ to use $t$ as the time for your canonical frame, but this is a global choice; you cannot do this on a per-query basis. However, with $E$’s Eulerian framing, we can extract a different point’s trajectory from the entirely different range $t^{\dagger}$ to $t^{\ddagger}$ (i.e. $E(x_{t^{\dagger}},y_{t^{\dagger}},z_{t^{\dagger}},t^{\dagger},t^{\ddagger})$) without concern for a canonical frame definition. It need not even be the case that $t<t^{\prime}$; indeed, this extraction works even if $t>t^{\prime}$, i.e. extracting the backwards trajectory through time.
D.3 Euler Integration to approximately solve the ODE
If $F$ is of arbitrary form and we want to compute the concrete values of $L$, we cannot exactly compute the continuous integral from 00 to $t$; we must approximate this with finite differences. Thus, we split the time range 00 to $t$ into $k$ steps, where each step is of size $\frac{t}{k}$. Thus, we can again define $L$ via recursion, but this time explicitly.
$L(x_{0},y_{0},z_{0},0)$ $=\left({x_{0},y_{0},z_{0}}\right)$ (12)
$L(x_{0},y_{0},z_{0},\tau+\frac{t}{k})$ $\approx L(x_{0},y_{0},z_{0},\tau)+\frac{t}{k}\cdot F(x_{\tau},y_{\tau},z_{\tau},\tau),$ (13)
or directly without recursion,
$L(x_{0},y_{0},z_{0},t)\approx\left({x_{0},y_{0},z_{0}}\right)+\sum_{n=1}^{k}\frac{t}{k}\cdot F(x_{n\frac{t}{k}},y_{n\frac{t}{k}},z_{n\frac{t}{k}},n\frac{t}{k})$ (14)
This finite difference solving approach is Euler integration.
D.4 Estimating the flow field with EulerFlow’s neural prior
For a given scene, we do not have access to $L$ or $F$ directly; these are are the true functions that uniquely characterize the underlying motion of the scene that we are trying to estimate. For EulerFlow, we represent our estimate of the scene’s flow field $F$ with a neural prior, $\theta$, i.e.
$F(x,y,z,t)\approx\theta(x,y,z,t)$ (15)
and thus
$L(x_{0},y_{0},z_{0},t)\approx\left({x_{0},y_{0},z_{0}}\right)+\sum_{n=1}^{k}\frac{t}{k}\cdot\theta{}(x_{n\frac{t}{k}},y_{n\frac{t}{k}},z_{n\frac{t}{k}},n\frac{t}{k})$ (16)
and, using the arbitrary start and end definition from Appendix D.2, with $k$ steps from the range $t$ to $t^{\prime}$ and $\delta=\frac{t^{\prime}-t}{k}$
$E(x_{t},y_{t},z_{t},t,t^{\prime})\approx E_{\theta}(x_{t},y_{t},z_{t},t,t^{\prime})=\left({x_{t},y_{t},z_{t}}\right)+\sum_{n=1}^{k}\delta\cdot\theta{}(x_{n\delta+t},y_{n\delta+t},z_{n\delta+t},n\delta+t)$ (17)
This formulation makes EulerFlow highly flexible, enabling optimization of $\theta$’s estimate of $F$ with objectives that take either an Eulerian view (directly on $\theta$ via Equation 15) or a Lagrangian view (on point rollouts for arbitrary start and end ranges via Equation 17).